OpenStudy (anonymous):

simplify: (-2x^2y^3z)^3=

OpenStudy (turingtest):

use the two properties \[(xy)^n=x^ny^n\]and\[(x^a)^b=x^{ab}\]

OpenStudy (anonymous):

huh? does that mean that both are the answer?

OpenStudy (turingtest):

no, I'm saying use those general rules to find your answer

OpenStudy (anonymous):

help??? any1 else whos in the room

OpenStudy (turingtest):

awww, I've been slighted :( good luck!

OpenStudy (anonymous):

Information given by "Turing Test" is enough to solve the question.

OpenStudy (anonymous):

(-2x^2y^3z)^3= -2^3*x^2^3*y^3^3*z^3

OpenStudy (anonymous):

i got it thnx

OpenStudy (across):

Make an effort to understand what TuringTest has said, and I promise you that you'll never have to ask us a question regarding this particular subject.

OpenStudy (anonymous):

im srry, im sick and not thinking to weel

OpenStudy (anonymous):

well

OpenStudy (anonymous):

and i got it